# CLAMP: Crowdsourcing a LArge-scale in-the-wild haptic dataset with an open-source device for Multimodal robot Perception

### v1

@@ 5 The CLAMP Model
The CLAMP model is a multimodal robot perception model that learns object material from visual and multimodal haptic data collected in-the-wild. The CLAMP model fuses a haptic encoder and a visual encoder by training on the CLAMP dataset.
@@
Haptic Encoder. Our haptic encoder processes time‐series haptic input through six InceptionTime [51] blocks, followed by a global average pooling layer, and finally a three‐layer MLP as the classification head. We split the CLAMP dataset sample-wise, train the encoder over three random seeds, and use the worst-performing seed for comparisons (see appendix for details).
Visual Encoder. We use GPT-4o [50] as a low-dimensional visual encoder. We take a two-step prompting approach: we first pass the object image and prompt to identify the object instance, and then use the instance prediction to prompt for a material class from the list of 14 materials. We extract per-material probabilities and normalize to get a class distribution (see appendix for details).
@@
The CLAMP Model: Combining Haptics and Vision. The CLAMP model uses a 2-layer MLP to fuse the outputs of the haptic and visual encoders. To train the combined model, we freeze the haptic encoder and minimize a composite loss function: $\mathcal{L}=\mathcal{L}_{WCE}+\lambda_{KL}\cdot\mathcal{L}_{KL}(\mathcal{V}||\mathcal{P})$ , where $\mathcal{L}_{WCE}$ is a weighted cross-entropy loss, $\lambda_{KL}$ is a weighting factor for KL divergence, and $\mathcal{L}_{KL}(\mathcal{V}||\mathcal{P})$ is a KL-divergence loss term with the vision probabilities $\mathcal{V}$ and the visuo-haptic model probabilities $\mathcal{P}$ [^1] We observe empirically that treating the visuo-haptic probabilities as the target distribution performs better than treating it as the source distribution..
@@
Comparison to Pretrained Vision Foundation Models. We next compare the CLAMP model to GPT-4o [50], CLIP [52], and PG-VLM [42], a physically-grounded vision-language model with state-of-the-art performance on material classification (among other tasks). CLIP is finetuned on CLAMP (see appendix for details). For each model, we consider three variations: raw images with no processing, images cropped to the object instance using Grounding Dino [53], and images segmented by Grounded SAM [54]. Table 2 reports test-set accuracy and normalized Matthews correlation coefficient (nMCC) on held-out data from the CLAMP dataset. We see that the CLAMP model consistently achieves the best performance. This suggests that the background knowledge of the pretrained models is not enough to make up for their lack of haptic input.
@@
Haptic Modality Ablation Studies. To further analyze the contributions of haptics, we ablate each of the five haptic modalities and compare classification performance to the full haptic encoder (without vision). See appendix for ablation details. In Table 2, we see that the full encoder performs best, suggesting the importance of each modality. Different sensing modalities provide complementary information towards material recognition. For example, active thermal sensing helps distinguish materials with different thermal effusivities [41], while force sensing distinguishes between materials in rigid and compliant objects [15]. Finally, we see that the CLAMP model performs better than the haptic encoder alone, confirming that both vision and haptics are necessary.
@@
Alternative Haptic Encoders. We next consider alternative haptic encoders. We first compare our InceptionTime encoder to a Random Forest encoder used in previous work on haptic classification [55], finding InceptionTime to obtain better performance (Table 2).
@@
Testing on Held-out Objects. We train a new CLAMP model on a dataset split into training, validation, and test sets by objects rather than by contact instances. The CLAMP model outperforms the vision-only baseline on the new test set, suggesting that our model can infer object material when tested on novel objects.
Transferring the Haptic Encoder to Compliance Recognition. We next consider whether our haptic encoder, which was trained for material classification, can also be used for object compliance recognition. We train a new classification head on the frozen haptic encoder using object compliance recognition labels (soft and hard) from the CLAMP dataset. The haptic encoder achieves an accuracy of 0.88 on the test set and an nMCC score of 0.869, suggesting that our haptic encoder learns representations that are applicable beyond material recognition.
Analyzing CLAMP Model Uncertainty. Uncertainty quantification is important for robot manipulation tasks to reduce failures  [56]. To assess whether CLAMP model predictions carry useful uncertainty signals, we mark a prediction as uncertain if the corresponding softmax probability is less than a threshold $p_{1}$, or if the difference between the largest and next largest softmax value is less than a threshold $p_{2}$. Filtering predictions increases the model test accuracy from 0.73 to 0.82, and the nMCC from 0.85 to 0.90, while retaining $75\%$ of predictions. We outline the values of $p_{1}$ and $p_{2}$ in the appendix.
Haptic Models Method Acc. nMCC InceptionTime 0.59 $\pm$ 0.01 0.76 $\pm$ 0.006 Random Forest 0.52 $\pm$ 0.001 0.71 $\pm$ 0.001 Modality-specific Ablation Absent Modality Acc. nMCC None 0.58 0.76 Active Therm. 0.49 0.7 Passive Therm. 0.46 0.68 Force 0.22 0.57 Vibration 0.57 0.75 Proprioception 0.54 0.74 Vision Models Method Acc. nMCC GPT-4o (raw) 0.65 0.80 GPT-4o (segmented) 0.59 0.77 GPT-4o (cropped) 0.64 0.80 CLIP (raw) 0.58 0.75 CLIP (segmented) 0.56 0.74 CLIP (cropped) 0.57 0.74 PG-VLM (raw) 0.39 0.67 PG-VLM (segmented) 0.44 0.69 PG-VLM (cropped) 0.58 0.76 Visuo-haptic models Method Acc. nMCC CLAMP model 0.87 0.93 CLIP + Haptic encoder 0.79 0.89 Evaluation on held-out objects Method Acc. nMCC CLAMP model 0.73 0.85 GPT only 0.70 0.83
Table 2: We report the performance of material recognition models on the CLAMP dataset.
@@ 6 Demonstrating the CLAMP Model on Real Robots
    Franka with CLAMP Device: We mount a shortened CLAMP device on the end effector of the Franka Panda. The device is actuated by tensioning its central steel strips, which are connected to a pulley system driven by a pair of interlocked Dynamixel XC330-M288-T motors. A 3D printed housing, mounted to the Franka, contains the motors and Pico, which actuates the device and monitors the sensor suite.
@@
The dataset size for the second and third embodiments is smaller than that for the first one, as the narrow grasping width prevents grasps for several household objects. We post-process and featurize the collected haptic data following the steps described in Section 4. However, for embodiments with a parallel jaw gripper, we modify the impedance feature calculation to use linear acceleration instead of angular velocity. Refer to appendix for details.
We finetune a pretrained CLAMP model with images collected offline with a CLAMP device and varying amounts (7%, 15% and 30% data) of embodiment-specific haptic data and the images of the corresponding objects collected offline with a CLAMP device. Table 3 shows that increasing the amount of finetuning data improves model performance, and that finetuned models for all three embodiments perform better than vision-based material predictions with $15\%$ of finetuning data.
@@
Leveraging multimodal perception for real-world manipulation. We finetune a CLAMP model with 85% of the finetuning dataset for the “Franka with Parallel Jaw” embodiment equipped with a wrist camera and deploy it in three real-world scenarios.
@@
    Sorting recyclable from non-recyclable waste: We conduct an experiment where the robot disposes 10 unseen objects into “Trash” or “Recycle” bins based on material predictions from the CLAMP model (sorting rules detailed in the appendix). Additionally, we incorporate thresholding for unknown and uncertain predictions to reduce false classifications and thus improve robustness. We find that the robot disposes objects in the correct bin 9/10 times per trial, while generating unknown predictions 0.66/10 times per trial over three trials.
@@
    Retrieving objects from a cluttered bag: We next conduct an experiment where the robot interacts with objects in a cloth bag laid flat on the table, and must retrieve the object only if it is metallic (aluminium, brass, or steel). The bag contains three unseen objects, one of which is metallic. We program actions for the robot to capture an image, reach into the bag, grasp each object, and retrieve it if it is identified by the CLAMP model as metallic. While the robot’s actions are determined by CLAMP model, we also log predictions from the vision-only model. The finetuned CLAMP model retrieved the metallic object in 6 out of 13 trials, while the vision model failed to identify the metallic object in any trial.
@@
    Sorting ripe from overripe bananas. In our final experiment, the robot must reason about object compliance to distinguish ripe from overripe bananas. The robot interacts with 4 bananas in a fruit basket (2 ripe, 2 overripe) and must only place ripe bananas on a plate. Our haptic encoder finetuned on embodiment-specific data achieves an accuracy of 0.833 and an nMCC of 0.86.
@@ 7 Limitations and Future Work
In this work, we presented the CLAMP device, dataset, and model, which collectively represent a significant step towards scaling visuo-haptic perception for robotics. We conclude with a discussion of limitations in the current framework and opportunities for future work.
CLAMP Device Design Constraints. The CLAMP device was optimized for weight, cost and portability. This lightweight design came with trade-offs. For example, we opted against using a more powerful microprocessor like the Raspberry Pi 4, which would have enabled higher bandwidth data collection, such as video streams, high-frequency contact microphone data, and vision-based tactile sensing. In addition, the current suction cup design allows for agile handling, but its low surface area and convex sensor bed sometimes hinder consistent contact with objects that have curved surfaces.
@@ Acknowledgments
This work was partly funded by National Science Foundation IIS #2132846, and CAREER #2238792. We thank Ruolin Ye, Rishabh Madan, Qi Chen, and other members of the EmPRISE Lab for feedback on the CLAMP device and for assistance in data collection.
@@ A.1 Hardware details
Force sensing. The CLAMP device uses an Interlink UX 402 force-sensing resistor (FSR) that can measure forces up to 150N. We calibrated the force-sensing resistor with an FX29 load cell. The resulting calibration curve is described by an exponential function of the voltage, achieving an R2 = 0.980. While not as accurate as load cell sensors or MEMS force sensors over long-term cyclic use, force-sensing resistors offer significant advantages in terms of resilience and cost.
@@
Proprioceptive sensing. An important aspect of haptics is that sensing depends on action. The contact forces generated while grasping an object depend on the velocity and the angle at which the object is grasped. To address this, the CLAMP device has two 6-axis MPU6050 IMUs. The axes of the two IMUs are oriented such that the Y-axis of IMU-2 aligns with the Z-axis of IMU-1, while the x-axis of IMU-2 is transformed at an angle of $-25\degree$ from that of IMU-1, about the Z-axis of IMU-1.
@@ B.1 Data processing
    We convert raw voltage readings from the active and passive thermal sensors to °C using the resistor values provided in the datasheet.  [^2] Despite access to a curve fit obtained from calibration, we do not convert raw voltage readings to readings in newtorns for the FSR. We train our model on uncalibrated force values. We verified that a material recognition model trained on calibrated force values performs equally as well as one trained on uncalibrated force values.
@@
    We create the nine features from the smoothed sensor data. Specifically,
@@
      $Z(t)$ =$=\begin{cases}\dfrac{F^{\prime}(t)}{\omega(t)}\;\;\;&\text{if }\omega(t)\geq\delta\\[5.0pt] 0\;\;\;&\text{if }\omega(t)<\delta\end{cases}$
      where $F^{\prime}(t)$ denotes the force difference at time $t$, $\omega(t)$ denotes the angular velocity feature, and $\delta$ denotes a threshold for angular velocity, which we fix as $3^{\circ}/s$ for our experiments. While impedance is typically computed via linear velocity, we use angular velocity because our IMU-based proprioception is more accurate in this dimension. We fix a lower bound to exclude spurious values of high impedance that we observe at low angular velocities. This often happens when users attempt to change the grasp contact point without releasing their grasp, resulting in a sudden change in force measurement from the FSR.
@@ B.2 Synchronizing contact
  • Contact on the right cup is determined by the force sensor readings. The cup is assumed to not be in contact at the initial time step. Contact is detected when the cup was not in contact on the previous timestep, and the force signal exceeds a certain threshold. Contact is released when the sensor was in contact and the force falls below the same threshold. We set the threshold as 0.01V.
@@ B.3 Filtering dataset for model learning
  • Contact instances where initial active temperature is less than 51 °$\degree C$. Prior work has shown that material recognition is challenging under varying initial conditions [18].
@@
  • Contact instances with maximum angular velocity below a threshold (chosen as 1°$\degree/s$, indicating a very slow grasp.
  • Contact instances with objects that have heterogenous surfaces.

### v2

@@ 5 The CLAMP Model
The CLAMP model is a multimodal perception model that recognizes object material from visual and multimodal in-the-wild haptic data. The model fuses a haptic encoder trained on the CLAMP dataset with a pretrained visual encoder.
@@
Haptic Encoder. Our haptic encoder is a six-block InceptionTime model  [31], followed by a global average pooling layer, and finally a three‐layer MLP as the classification head. We split the CLAMP dataset sample-wise, train the encoder over three random seeds, and use the worst-performing model for comparisons. See Appendix C.1 for more details.
Visual Encoder. We use GPT-4o [42] as a low-dimensional visual encoder. Given an input image, audio transcription, and in-context examples, it outputs per-token probabilities, which we filter and normalize into a categorical distribution over material classes. See Appendix C.2 for more details.
@@
The CLAMP Model: Combining Haptics and Vision. The CLAMP model uses a 2-layer MLP to fuse the outputs of the haptic and visual encoders. To train the model, we freeze the pretrained haptic encoder and minimize a loss function: $\mathcal{L}=\mathcal{L}_{WCE}+\lambda_{KL}\cdot\mathcal{L}_{KL}(\mathcal{V}||\mathcal{P})$ , where $\mathcal{L}_{WCE}$ is a weighted cross-entropy loss, $\mathcal{L}_{KL}(\mathcal{V}||\mathcal{P})$ is a KL-divergence loss term with the vision probabilities $\mathcal{V}$ and the visuo-haptic model probabilities $\mathcal{P}$, and $\lambda_{KL}$ is the weight for KL loss (Appendix C.1).
@@
Comparison to Pretrained Vision Foundation Models. We compare the CLAMP model on material classification performance to GPT-4o [42], a grounded vision-language model (PG-VLM [29]), and a CLIP [43] model finetuned on our dataset. For each model, we consider three inputs: raw images, images cropped to the object instance using Grounding DINO [39], and images segmented by Grounded SAM [47]. We compare the CLAMP model with the three vision-only models, across all inputs, in Table 2, which reports test-set accuracy and normalized Matthews correlation coefficient (nMCC). The CLAMP model achieves the best performance, suggesting that the background knowledge of the pretrained models cannot compensate for their lack of haptic input.
@@
Haptic Modality Ablation Studies. We compare classification performance of the haptic encoder to ablated encoders trained on all but one haptic modality. In Table 2, we see that the encoder trained on all modalities performs best, suggesting the importance of each modality. All five modalities provide complementary cues about object properties. Force conditioned on proprioception distinguishes soft objects from hard ones [7], vibration sensing captures high-frequency contact cues in hard objects, and active/passive thermal sensing separates objects of similar compliance by thermal conductivity and temperature [4]. Finally, we see that the CLAMP model performs better than the haptic encoder alone, confirming that both vision and haptics are necessary for material recognition.
@@
Alternative Haptic Encoders. We consider alternative network architectures for the haptic encoder. We first compare our InceptionTime encoder to a Random Forest encoder used in previous work on haptic classification [23], finding InceptionTime to obtain better performance (Table 2).
@@
Testing on Held-out Objects. We train a new CLAMP model on a dataset split into training, validation, and test sets by objects rather than by contact instances. The CLAMP model outperforms the vision-only baseline on the new test set, suggesting that our model can infer object material when tested on novel objects. The CLAMP model’s 3% accuracy improvement over the vision-only baseline is statistically significant with $p\>\textless\>0.05$ (using the Wilcoxon-Signed Rank test).
Transferring the Haptic Encoder to Compliance Recognition. We show that our haptic encoder, which was trained for material classification, can be transferred to recognize object compliance. We train a new classification head on the frozen haptic encoder using object compliance recognition labels (soft and hard) from the CLAMP dataset. The haptic encoder achieves an accuracy of 0.88 on the test set and an nMCC score of 0.869. This result suggests that our haptic encoder learns representations that are applicable beyond material recognition.
Analyzing CLAMP Model Uncertainty. Uncertainty quantification is important for robot manipulation tasks to reduce failures  [46]. To assess whether CLAMP model predictions carry uncertainty signals, we mark a prediction as uncertain if the corresponding softmax value is less than a threshold $p_{1}$, or if the difference between the two largest softmax values is less than a threshold $p_{2}$. Filtering predictions increases the model test accuracy from 0.73 to 0.82, and the nMCC from 0.85 to 0.90, while retaining $75\%$ of predictions. Refer to Appendix C.1 for values of $p_{1}$ and $p_{2}$.
Haptic Models Method Acc. nMCC InceptionTime 0.59 $\pm$ 0.01 0.76 $\pm$ 0.006 Random Forest 0.52 $\pm$ 0.001 0.71 $\pm$ 0.001 Modality-specific Ablation Absent Modality Acc. nMCC None 0.58 0.76 Active Therm. 0.49 0.7 Passive Therm. 0.46 0.68 Force 0.22 0.57 Vibration 0.57 0.75 Proprioception 0.54 0.74 Vision Models Method Acc. nMCC GPT-4o (raw) 0.65 0.80 GPT-4o (segmented) 0.59 0.77 GPT-4o (cropped) 0.64 0.80 CLIP (raw) 0.58 0.75 CLIP (segmented) 0.56 0.74 CLIP (cropped) 0.57 0.74 PG-VLM (raw) 0.39 0.67 PG-VLM (segmented) 0.44 0.69 PG-VLM (cropped) 0.58 0.76 Visuo-haptic models Method Acc. nMCC CLAMP model 0.87 0.93 CLIP + Haptic encoder 0.79 0.89 Evaluation on held-out objects Method Acc. nMCC CLAMP model 0.73 0.85 GPT-only 0.70 0.83
Table 2: We report the performance of material recognition models on the CLAMP dataset. Note that for “Haptic Models”, metrics are reported as mean $\pm$ stddev.
@@ 6 Demonstrating the CLAMP Model on Real Robots
    Franka with CLAMP Device: We mount a shortened CLAMP device on the end effector of the Franka Panda. The device is actuated by tensioning its central steel strips, which are connected to a pulley system driven by a pair of interlocked Dynamixel XC330-M288-T motors. A 3D printed housing contains the motors and the Pico.
@@
The narrow grasping width for the two parallel jaw embodiments prevents grasps for several household objects, resulting in a smaller dataset for these embodiments compared to that for the first one. We post-process and featurize the collected haptic data following the steps described in Section 4. However, for embodiments with a parallel jaw gripper, we modify the impedance feature calculation to use linear acceleration instead of angular velocity. Refer to Appendix D.1 for details.
We finetune a pretrained CLAMP model with varying amounts (7%, 15%, and 30% data) of embodiment-specific haptic data and offline images collected with a CLAMP device. Table 3 shows that using more finetuning data improves model performance, and that finetuned models for all three embodiments perform better than vision-based material predictions with $15\%$ of finetuning data.
@@
Leveraging multimodal perception for real-world manipulation. We finetune a CLAMP model with 85% of the finetuning dataset for the “Franka with Parallel Jaw” embodiment equipped with a wrist camera, and deploy it in three real-world scenarios.
@@
    Sorting recyclable waste from trash: In this setting, the robot disposes of 10 objects not seen in the finetuning dataset, into “Trash” or “Recycle” bins based on material predictions from the CLAMP model (sorting rules detailed in Appendix D.3), with an “I’m not sure” box for uncertain predictions. The robot disposes of objects in the correct bin 9/10 times per trial, while generating unknown predictions 0.66/10 times per trial, over three trials. This shows that our model can accurately recognize material for unseen objects in cluttered environments.
@@
    Retrieving objects from a cluttered bag: In this setting, the robot interacts with objects inside a cloth bag, and must retrieve the object only if it is metallic. The bag contains three unseen objects, one of which is metallic. We also log predictions from the vision-only model for comparison. The finetuned CLAMP model retrieved the metallic object in 6 out of 13 trials, while the vision model does not find a metallic object in any of the 13 trials. We thus show that haptics provides key cues in visually challenging situations. See Appendix D.4 for more details.
@@
    Sorting ripe from overripe bananas: In this setting, the robot interacts with four bananas (two of which are overripe), predicts its compliance, and retrieves only ripe bananas. Our haptic encoder transferred to compliance prediction and finetuned on robot data achieves an accuracy of 0.833 and an nMCC of 0.86. This shows that our model can differentiate between visually ambiguous objects of the same instance, based on object compliance.
@@ 7 Limitations and Future Work
In this work, we present the CLAMP device, dataset, and model, which collectively represent a significant step towards scaling visuo-haptic perception for robotics. We conclude with a discussion of limitations in the current framework and opportunities for future work.
CLAMP Device Design Constraints. The CLAMP device was optimized for weight, cost and portability. This lightweight design came with trade-offs. For example, we opted against using a more powerful microprocessor like the Raspberry Pi 4, which would have enabled higher bandwidth data collection such as video streams, high-frequency contact microphone data, and vision-based tactile sensing. In addition, the current suction cup design allows for agile handling, but its low surface area and convex sensor bed can hinder consistent contact with objects with curved surfaces.
@@ Acknowledgments
This work was partly funded by National Science Foundation IIS #2132846, and CAREER #2238792. We thank Ruolin Ye, Rishabh Madan, Qi Chen, and other members of the EmPRISE Lab for feedback on the CLAMP device, assistance with figures in the manuscript, and assistance in data collection.
@@ A.1 Hardware details
Force sensing. The CLAMP device uses an Interlink UX 402 force-sensing resistor (FSR) that can measure forces up to 150 N. We calibrate the force-sensing resistor with an FX29 load cell. The resulting calibration curve is described by an exponential function of the voltage, achieving an R2 = 0.980. While not as accurate as load cell sensors or MEMS force sensors over long-term cyclic use, force-sensing resistors offer significant advantages in terms of resilience and cost.
@@
Proprioceptive sensing. An important aspect of haptics is that sensing depends on action. The contact forces generated while grasping an object depend on the velocity and the angle at which the object is grasped. To address this, the CLAMP device has two 6-axis MPU6050 IMUs. The axes of the two IMUs are oriented such that the Y-axis of IMU-2 aligns with the Z-axis of IMU-1, while the x-axis of IMU-2 is transformed at an angle of $25\degree$ from that of IMU-1, about the Z-axis of IMU-1.
@@ B.1 Data processing
    We convert raw voltage readings from the active and passive thermal sensors to °C using the resistor values provided in the datasheet.  [^1] Despite access to a curve fit obtained from calibration, we do not convert raw voltage readings to readings in Newtons for the FSR. We train our model on uncalibrated force values. We verified that a material recognition model trained on calibrated force values performs equally as well as one trained on uncalibrated force values.
@@
    We extract a total of nine features from all sensor data streams. For the thermal sensors, the force sensor, and the contact microphone, we use the smoothed readings directly. For the IMU data, we compute the angular velocity of the end-effector IMU relative to the IMU in the gripper body. This results in five primary features. We then create four additional features: differences in active thermal, passive thermal, and force readings between consecutive timesteps, and an impedance feature. Note that:
@@
      $Z(t)$ $=\begin{cases}\dfrac{F^{\prime}(t)}{\omega(t)}\;\;\;&\text{if }\omega(t)\geq\delta\\[5.0pt] 0\;\;\;&\text{if }\omega(t)<\delta\end{cases}$
      where $F^{\prime}(t)$ denotes the force difference at time $t$, $\omega(t)$ denotes the angular velocity feature, and $\delta$ denotes a threshold for angular velocity, which we fix as $3^{\circ}/s$ for our experiments. We find that this impedance feature reliably indicates contact and shows distinct responses for contact with rigid and compliant objects. While impedance is typically computed via linear velocity, we use angular velocity because our IMU-based proprioception is more accurate in this dimension. We fix a lower bound to exclude spurious values of high impedance that we observe at low angular velocities. This often happens when users attempt to change the grasp contact point without releasing their grasp, resulting in a sudden change in force measurement from the FSR.
@@ B.2 Synchronizing contact
  • Contact on the right cup is determined by the force sensor readings. The cup is assumed to not be in contact at the initial timestep. Contact is detected when the cup was not in contact on the previous timestep, and the force signal exceeds a certain threshold. Contact is released when the sensor was in contact and the force falls below the same threshold. We set the threshold as 0.1V.
@@ B.3 Filtering dataset for model learning
  • Contact instances where initial active temperature is less than 51 $\degree C$. Prior work has shown that material recognition is challenging under varying initial conditions [11].
@@
[…]
  • Contact instances with objects that have heterogeneous surfaces.